OpenStudy (anonymous):

What equation would I use to model this situation? A Broadway musical was attended by a total of 2,838 people, including men, women, and children. The number of women who attended the event was 5 more than 3 times the number of children, while the number of men was 3 less than 4 times the number of women. Write an equation to model this situation. Then, use the equation to find the number of men, women, and children who attended the musical.

OpenStudy (jdoe0001):

@A-HUGE-EGG any ideas yet?

OpenStudy (anonymous):

w=5+(3*c) m=(4*w)-3 So you want to find the formula for children 3c+5=w 3c=w-5 c=(w-5)/3 plug that into women w=5+(3*(w-5/3)) w=5+((3w-15)/3) w=5+w-15 HAHA well I got this far but now it cancels out w

OpenStudy (anonymous):

Okay I think I got it If you take the formula for W=5+3c and M=4w-3 if you plug w into m you get m=4(5+3c)-3 simplify and you get m=20+12c-3 m=17+12c Now what you can do is solve for c in that last formula and you get 12c=m-17 c=(m-17)/12 So if you look at everything we have one formula for men, one for women, but two for children: m=4w-3 w=5+3c c=(w-5)/3 and c=(m-17)/12 so let us use the second children formula and plug it into women, then we get w=5+3((m-17)/12) w=5+((3m-51)/12) plug men into that w=5+(3*(4w-3)-51)/12 you get w=5+(12w+153)/12 w=5+w+12.75 And that is where I give up.... Sorry

OpenStudy (anonymous):

I am determined to figure this out!!!!

OpenStudy (anonymous):

I got more people to help me! You are not alone!